# Conceptualization, Methodology, Writing – original draft, Software, Writing – review & editing

### v2

@@ 5 Modeling estimation result
This study develops a series of Bayesian hierarchical UGEV models to estimate COR from extreme 2D-TTC near-miss events observed along Biscayne Boulevard in Miami. Extreme near-misses are identified using the BM approach, wherein the minimum 2D-TTC within each temporal block is treated as the most critical interaction. The theoretical foundations and implementation of this sampling methodology are detailed in prior research (2025b). Each V–V and V–I interaction is treated as an independent block. Because extreme events are sparse at individual sites, direct location-specific estimation can yield unstable posteriors with wide uncertainty intervals. To address this, a grouping strategy was adopted, pooling samples across functionally similar sites (intersections, segments) while still preserving local heterogeneity. This clustering approach improves estimation stability and aligns with prior network-level applications of UGEV models in traffic safety (Zheng et al., 2014; Fu and Sayed, 2021c).
@@
The primary analytical framework is the HBSGRP–UGEV model, while a simpler fixed-effect hierarchical model (HBSFP) was estimated as a benchmark. Accordingly, two models were developed for each near-miss type: (1) HBSFP and (2) HBSGRP. The grouped random parameter specification allows selected coefficients to vary across spatial clusters, thereby capturing latent differences linked to roadway geometry and interaction dynamics. Model performance was evaluated using the DIC (Spiegelhalter et al., 2002; El-Basyouny and Sayed, 2012b, a), where lower values indicate stronger predictive performance (Zheng et al., 2019).
@@
Estimation was conducted using MultiBUGS, a parallelized extension of OpenBUGS tailored for high-dimensional Bayesian inference. Its support for block updating and parallel execution makes it well-suited for capturing spatial dependence, random effects, and extreme value behavior within complex data structures. Parameters were estimated by Markov Chain Monte Carlo (MCMC) simulation, two parallel MCMC chains of 50000 iterations were run, with the first 20000 discarded as burn-in and the remaining 30000 retained for posterior inference. Convergence was verified both visually (trace plots) and quantitatively (Brooks–Gelman–Rubin diagnostic), with all monitored parameters achieving BGR values below 1.1, indicating satisfactory convergence (Gelman and Rubin, 1992; Brooks and Gelman, 1998; El-Basyouny and Sayed, 2009). This estimation procedure yielded robust posterior summaries of means, standard deviations, and 95% credible intervals, even under correlation structures induced by spatial clustering and random effects.
@@
The study incorporated a comprehensive set of covariates into the location ($\mu$) and scale ($\sigma$) parameters of the non-stationary UGEV distribution to reflect contextual and behavioral variability in COR. For both V–V and V–I interactions, dynamic covariates included relative speed, distance, acceleration, deceleration, headway difference, steering angle difference, jerk and traffic volume, factors capturing key dimensions of vehicle dynamics and near-miss severity. Geometric and infrastructure related covariates were also included: lane number, lane width, driveway intensity, median type, and lane position for segment models; and a categorical indicator of turning movement for intersection models.
@@
All covariates were specified in both $\mu$ and $\sigma$ components, enabling the models to capture variation in both the central tendency and dispersion of extreme 2D-TTC. In contrast, the shape parameter ($\xi$) was held constant across all locations. This choice is deliberate: $\xi$ governs tail heaviness in the UGEV distribution but is notoriously difficult to estimate from sparse BM samples. Allowing $\xi$ to vary with covariates often induces instability and overfitting (Coles et al., 2001; Cooley et al., 2006). Fixing $\xi$ improves convergence and yields more interpretable posterior distributions.
@@ 5.1 Posterior results and model performance
Tables 5 through 8 report the posterior summaries for both HBSFP and HBSGRP models, estimated separately for V–V and V–I near-misses at intersections and segments. Across all cases, the HBSGRP specification consistently outperformed the HBSFP baseline, confirming its ability to capture unobserved heterogeneity and locatione-level variability. The DIC values for the HBSGRP models were 508.9 (V–V) and 3179.6 (V–I) at intersections, and 4189.0 (V–V) and 6550.0 (V–I) at segments. These correspond to relative reductions in DIC of 3.41% and 0.54% at intersections, and 7.49% and 3.11% at segments, compared with their HBSFP counterparts. Such reductions exceed the widely accepted threshold of five DIC units, indicating statistically meaningful improvements in fit (El-Basyouny and Sayed, 2012b, a). To further assess model robustness, WAIC and LOOIC were also computed. Both measures produced rankings consistent with DIC, reinforcing the superior performance of the HBSGRP structure. By allowing selected coefficients to vary across spatial groups, the HBSGRP models capture latent heterogeneity in both the central tendency and dispersion of near-miss severity, which fixed-effect structures fail to represent. Collectively, these results demonstrate that grouped random parameter models provide a more flexible and accurate framework for modeling COR dynamics in complex urban traffic environments.
@@ 5.1.1 Intersection-level
The intersection-level HBSGRP–UGEV model was designed to capture the heterogeneity of extreme 2D-TTC distributions arising from V-V interactions at signalized locations. In the model specification stage, several covariates were tested as random effects; however, most roadway geometry variables exhibited negligible variation across sites and did not improve posterior fit. To maintain parsimony and interpretability, random effects were therefore restricted to the intercept and two dynamic covariates (relative speed and relative distance), while other vehicle dynamics and geometric features were retained as fixed effects, following recommendations from (Anis et al., 2025a). This choice reflects theoretical expectations: vehicle dynamics are highly variable across intersections depending on signal phasing, approach alignment, and driver behavior, whereas roadway structure tends to exert a more stable influence. The resulting specification captures both site-specific dynamics and broader geometric constraints, group-specific random effects illustrated in Fig. 5.1.1 with posterior summaries reported in Table 5.
@@
The fixed-parameter benchmark (HBSFP) provides a baseline against which the grouped random specification can be compared. Consistent reductions in DIC, WAIC, and LOOIC confirm the superior fit of the HBSGRP model, underscoring the explanatory value of allowing relative speed and spacing effects to vary across intersection clusters. The direction and magnitude of these coefficients provide valuable insights into the mechanisms underlying COR. The random-parameter specification highlights the spatially varying influence of relative speed and relative distance on both the location and scale components of the HBSGRP distribution. For $\mu$, relative speed shows a positive effect across intersections, indicating that larger speed differentials between interacting vehicles lead to sharply reduced 2D-TTC values and more severe near-miss conditions. In contrast, relative distance is negatively associated with $\mu$, meaning that shorter headways systematically increase near-miss severity by limiting the buffer available for evasive action. For $\sigma$, both covariates are associated with reduced variability in near-miss severity. Relative speed has a negative coefficient, suggesting that high-speed interactions produce consistently critical outcomes, dangerous but less variable, because drivers have little opportunity to recover through braking or swerving. Similarly, relative distance exerts a negative effect on $\sigma$, reflecting that short headways yield uniformly critical near-miss with little dispersion in outcomes. Together, these results confirm the physical intuition that speed amplifies severity while distance compresses variability, producing deterministic near miss outcomes under constrained conditions. Fig. 5.1.1 further illustrate this spatial heterogeneity across intersection, while most locations show consistent trends, the I2 site exhibits unusually wide CI for the $\mu$ effects and CI exceed zero for the $\sigma$, across intercept, relative speed, and relative distance. This instability is directly attributable to I2’s small BM sample size, which weakens posterior precision and introduces greater uncertainty in estimated group-level effects. Despite this, the overall pattern across intersections remains robust: relative speed consistently intensifies near-miss severity, while shorter distances exacerbate risk and constrain outcome variability.
@@
Beyond the random effects, several covariates were modeled as fixed parameters due to their relatively stable contributions across intersections. Their interpretation reveals distinct roles of longitudinal dynamics, lateral control, maneuver type, and exposure effects in shaping near-miss severity and variability. Table 5 indicate that relative acceleration shows a significant negative effect on both the location and scale parameters. This suggests that acceleration is generally associated with smoother and less variable near-miss, consistent with situations where vehicles accelerate out of a green phase or clear an intersection without abrupt interaction. By contrast, relative deceleration exerts a positive effect on severity ($\mu$) but a negative effect on variability ($\sigma$). This indicates that abrupt braking sharply intensifies near-misses but also produces more uniform outcomes, reflecting the deterministic nature of emergency deceleration in high-risk encounters. Jerk, which captures the rate of change in acceleration, has the opposite pattern: a negative effect on $\mu$ but a positive effect on $\sigma$. Although jerk slightly reduces average near-miss severity (perhaps due to hesitation or driver corrections), it simultaneously increases unpredictability in outcomes, highlighting the unstable nature of rapid throttle/brake oscillations.
@@
Lateral vehicle dynamics also contribute meaningfully to near-miss risk estimation. Steering difference emerges as a significant contributor in the HBSGRP specification, showing a positive association with $\mu$ and a negative association with $\sigma$. This suggests that abrupt steering maneuvers intensify near-miss severity but lead to more consistent outcomes within constrained intersection geometries. Heading difference also shows a significant positively association with $\mu$ and a negatively with $\sigma$. Indicating that even small misalignments in trajectory lead to predictable patterns of near-miss outcomes, consistent with the rigid geometry of signalized approaches.
@@
The categorical indicator for left-turn movements is significant and has a positive effect on both $\mu$ and $\sigma$. This aligns with the well-documented complexity of left-turn phases, which involve interactions with opposing through traffic and diverse driver decision strategies. The result suggests that left turns not only elevate near-miss severity but also introduce greater variability in outcomes, reflecting the wide range of intersection designs and driver behaviors under turning conditions. Traffic volume is statistically insignificant in both $\mu$ and $\sigma$. Nonetheless, it was retained in the model to ensure conceptual completeness, since it represents exposure. The lack of effect implies that short-duration 2D-TTC near-miss events are governed more by instantaneous vehicle interactions than by aggregate traffic demand.
@@
The HBSGRP model for V–I near-misses specifies the intercept as a random parameter in both the location and scale components, capturing baseline heterogeneity in near-miss severity and variability across intersections. Relative speed (location) and relative distance (scale) were also treated as random effects (Fig. 5.1.1), though their 95% CI overlapped zero. They were retained on theoretical grounds, given their inclusion improve overall metrics. All other covariates were modeled as fixed effects, representing more stable behavioral and geometric influences. Fit statistics (Table 6) show modest but consistent improvements in DIC, WAIC, and LOOIC over the HBSFP benchmark, suggesting that the grouped random parameter structure adds nuance but yields smaller performance gains compared to the V–V case.
@@
In the location parameter, several interpretable effects emerge. Relative distance is strongly negative and significant, confirming that reduced clearance to the roadway edge sharply lowers 2D-TTC, elevating severity. Relative deceleration also carries a significant negative sign, which contrasts with its positive effect in V–V near-miss. This suggests that braking near infrastructure reflects proactive yielding or evasive action, thereby limiting severity rather than escalating it. Heading and steering differences are positive and significant, indicating that lateral deviations (e.g., turning toward or away from curbs) increase near-miss severity. Left-turn movement, however, shows a significant negative association, suggesting that left-turn trajectories may provide greater clearance to boundaries, reducing average severity compared to right turns. This finding requires careful interpretation, as it may partly reflect geometric factors such as wider turning radii or reduced pedestrian/boundary overlap in left-turn phasing. Relative speed, though included as a random effect, is not significant in the pooled analysis, highlighting that V–I risks are less sensitive to speed differentials than V–V encounters.
@@
The scale parameter results add further insight into variability. The negative random intercept suggests baseline differences across intersections, with some sites producing more predictable outcomes than others. Relative speed exhibits a positive and significant association with scale, implying that faster approaches to infrastructure yield greater variability in severity—contrasting with the V–V case, where speed reduced dispersion. This difference reflects the inherently diverse driver responses to boundary proximity: some brake aggressively, while others take wider trajectories or maintain speed. Relative acceleration is also positive and significant, reinforcing the idea that heterogeneous driver behavior during acceleration near boundaries widens outcome variability. In contrast, relative deceleration has a significant negative effect on scale, meaning that strong braking compresses the severity distribution into consistently critical outcomes. Jerk is similarly positive, confirming that abrupt longitudinal control amplifies variability. Lateral dynamics again exhibit distinct roles: heading and steering differences reduce variability, suggesting that once a driver commits to a lateral maneuver, outcomes become more predictable. Finally, left-turn movement increases variability despite lowering average severity, consistent with heterogeneous geometric layouts and driver strategies in left-turn contexts.
@@
The shape parameter remains negative and significant across intersections, supporting the Weibull domain for V–I near misses and confirming that these events have a bounded upper tail. This implies a physical or behavioral limit to how severe boundary near-misses can become, constrained by roadway geometry and driver response capacity. Spatial variation in the shape parameter underscores the role of intersection design and infrastructure placement in shaping the extremal properties of near-miss distributions. Overall, the intersection-level results suggest that V–I near-misses are driven more by clearance, braking behavior, and lateral maneuvers than by speed differentials. Unlike V–V near-misses, where high speeds produce deterministic severity, V–I cnear-misses are marked by greater variability, reflecting the complex interplay between infrastructure geometry, driver avoidance strategies, and intersection control.
5.2 Segment-level

### v3

@@ 5 Modeling estimation result
This study develops a series of Bayesian hierarchical UGEV models to estimate corridor-wide COR from extreme 2D-TTC near-miss events observed along Biscayne Boulevard in Miami. Extreme near-misses are identified using the BM approach, wherein the minimum 2D-TTC within each temporal block is treated as the most critical interaction. The theoretical basis and implementation of this sampling methodology are detailed in prior work (2025b). Each V–V and V–I interaction is treated as an independent block. Since extreme events are relatively sparse at individual sites, direct location-specific estimation often yields unstable posteriors with wide uncertainty intervals. To address this issue, a corridor-wide grouping strategy was adopted, pooling samples across functionally similar sites (intersections, segments) while preserving local heterogeneity.
@@
The HBSGRP–UGEV model is the primary analytical framework, with a simpler fixed-effect hierarchical model (HBSFP) also estimated for comparison. For each near-miss type (V–V and V–I), two models were therefore developed: (1) HBSFP and (2) HBSGRP. The grouped random parameter formulation in HBSGRP allows selected coefficients to vary across spatial clusters, thereby capturing latent heterogeneity associated with interaction dynamics and roadway characteristics. Model performance was evaluated using the DIC (Spiegelhalter et al., 2002; El-Basyouny and Sayed, 2012b, a), with lower values indicating superior predictive fit (Zheng et al., 2019).
@@
Estimation was conducted using MultiBUGS, a parallelized extension of OpenBUGS tailored for high-dimensional Bayesian inference. Its support for block updating and parallel execution makes it well-suited to modeling spatial dependence, random effects, and extreme value behavior within complex data structures. Parameters were estimated by Markov Chain Monte Carlo (MCMC) simulation; two parallel MCMC chains of 50000 iterations were run, with the first 20000 discarded as burn-in and the remaining 30000 retained for posterior inference. Convergence was assessed both visually (trace plots) and quantitatively (Brooks–Gelman–Rubin diagnostic), with all monitored parameters achieving BGR values below 1.1, confirming reliable convergence (Gelman and Rubin, 1992; Brooks and Gelman, 1998; El-Basyouny and Sayed, 2009). This procedure yielded robust posterior summaries of means, standard deviations, and 95% credible intervals, even under correlation structures induced by spatial clustering and random effects.
@@
The modeling framework incorporated a comprehensive set of covariates into the location ($\mu$) and scale ($\sigma$) parameters of the nonstationary UGEV distribution to capture contextual and behavioral variability in COR. For both V–V and V–I interactions, dynamic covariates included relative (speed, distance, acceleration, deceleration, headway, steering angle), jerk, and traffic volume, thereby capturing multiple dimensions of vehicle dynamics. Geometric and infrastructure related covariates were also included: lane number, lane width, driveway intensity, median type, and lane position of vehicle.
@@
All covariates were specified in both the $\mu$ and $\sigma$ components, enabling the models to capture variation in both the central tendency and dispersion of extreme 2D-TTC. In contrast, the shape parameter ($\xi$) was held constant across all locations. This decision reflects a tradeoff between flexibility and stability: although $\xi$ governs tail heaviness in the UGEV distribution, it is notoriously difficult to estimate from sparse BM samples. Allowing $\xi$ to vary with covariates often induces instability and overfitting (Coles et al., 2001; Cooley et al., 2006), whereas fixing it improves convergence and yields more interpretable posterior distributions.
@@ 5.1 Posterior results and model performance
Tables 5 through 8 report the posterior summaries for both HBSFP and HBSGRP models, estimated separately for V–V and V–I near-misses at intersections and segments. Across all cases, the HBSGRP structure consistently outperformed the HBSFP baseline model, confirming its capacity to capture unobserved heterogeneity and location-level variability. The DIC values for the HBSGRP models were 508.9 (V–V) and 3179.6 (V–I) at intersections, and 4189.0 (V–V) and 6550.0 (V–I) at segments. These correspond to relative reductions of 3.41% and 0.54% at intersections, and 7.49% and 3.11% at segments, compared with their HBSFP counterparts. Such improvements exceed the widely accepted threshold of five DIC units, indicating substantively better model fit (El-Basyouny and Sayed, 2012b, a). To further evaluate robustness, WAIC and LOOIC were also computed, both of which produced rankings consistent with DIC, thereby reinforcing the superior performance of the HBSGRP. By allowing selected coefficients to vary across spatial groups, the HBSGRP models capture latent heterogeneity in both the central tendency and dispersion, a capability absent in fixed-effect structures. Collectively, these results demonstrate that grouped random parameter models provide a more flexible and accurate framework for modeling COR dynamics in complex urban traffic environments.
@@ 5.1.1 Intersection-level
The intersection-level HBSGRP–UGEV model was designed to capture the heterogeneity of extreme 2D-TTC distributions from V-V interactions at signalized intersections. In the modeling stage, all covariates were considered as random effects; however, roadway characteristics and few vehicle dynamics were exhibited negligible variation across sites and did not improve posterior fit. To maintain parsimony and interpretability, random effects were restricted to the intercept and two dynamic covariates (relative speed and distance), while other covariates were retained as fixed effects, following recommendations from (Anis et al., 2025a). This choice reflects theoretical expectations: vehicle dynamics are highly variable across intersections depending on signal phasing, approach alignment, and driver behavior, whereas roadway structure tends to exert a more stable influence. The resulting specification captures both site-specific dynamics and broader geometric constraints, group-specific random effects illustrated in Fig. 5.1.1 with posterior summaries reported in Table 5.
@@
Consistent reductions in DIC, WAIC, and LOOIC confirm the superior fit of the HBSGRP model, highlighting the added value of allowing speed and spacing effects to vary across intersections. The direction and magnitude of these coefficients provide insights into the mechanisms underlying COR. Relative speed and distance emerge as the dominant dynamics, influencing both the location ($\mu$) and scale ($\sigma$) parameters. For $\mu$, relative speed has a positive effect, indicating that larger speed differentials between vehicles sharply reduce 2D-TTC and intensify near-miss severity. In contrast, relative distance shows a negative effect, meaning shorter headways systematically elevate severity by limiting available buffer space for evasive action. For $\sigma$, both covariates carry negative signs, implying that high speeds and short distances compress variability, yielding consistently critical outcomes with little room for driver recovery through braking or swerving. Taken together, these results confirm that speed amplifies severity while short headways exacerbate risk and reduce variability, producing deterministic near-miss outcomes under constrained conditions. Fig. 5.1.1 illustrates the spatial heterogeneity across intersections. While most locations show consistent trends, the I2 site exhibits unusually wide credible intervals for the $\mu$ effects and intervals exceeding zero for the $\sigma$ across the intercept, relative speed, and relative distance. This instability is attributable to I2’s small BM sample size, which weakens posterior precision and increases uncertainty in group-level effects. Despite localized uncertainty, the dominant intersection-level trend persists, with relative speed and distance exerting stable and significant effects on near-miss outcomes.
@@
Beyond random effects, several covariates were modeled as fixed parameters due to their stable contributions across intersections. These effects can be interpreted in terms of driver behavior. These factors highlight the roles of longitudinal dynamics, lateral control, maneuver type, and exposure in shaping near-miss outcomes. Table 5 shows that relative acceleration has a significant negative effect on both the location and scale parameters, suggesting smoother and less variable near misses when vehicles accelerate through green phases or clear intersections without abrupt interaction. By contrast, relative deceleration increases severity ($\mu$) but reduces variability ($\sigma$), capturing emergency braking events. These situations sharply intensify near misses but also produce uniform outcomes, since drivers in crisis have limited response options. Jerk, representing the rate of change in acceleration, shows the opposite pattern: a negative effect on $\mu$ but a positive effect on $\sigma$. This suggests that rapid throttle–brake oscillations, often signs of hesitation or overcorrection, slightly lower average severity but introduce greater unpredictability, underscoring the instability of abrupt corrective maneuvers.
@@
Lateral vehicle dynamics also play an important role in near-miss risk estimation. Steering difference is a significant factor in the HBSGRP model, showing a positive association with $\mu$ and a negative association with $\sigma$. This indicates that abrupt steering maneuvers intensify near-miss severity but produce more consistent outcomes within constrained intersection geometries. Heading difference shows a similar pattern, with a positive effect on $\mu$ and a negative effect on $\sigma$. This suggests that even small misalignments in vehicle trajectory elevate severity while reducing variability, reflecting the rigid geometric constraints of signalized approaches.
@@
The categorical indicator for left-turn movements is significant, with positive effects on both $\mu$ and $\sigma$. This finding is consistent with the well-documented complexity of left-turn phases, which involve interactions with opposing through traffic and diverse driver decision strategies. Left turns therefore not only elevate near-miss severity but also increase variability, reflecting differences in intersection design and driver behavior under turning conditions. By contrast, traffic volume is statistically insignificant for both parameters but was retained in the model to represent exposure. Its lack of effect suggests that short-duration extreme 2D-TTC events are driven more by instantaneous vehicle interactions than by aggregate demand.
@@
The HBSGRP model for V–I near misses specified the intercept as a random parameter in both the location and scale components, capturing baseline heterogeneity in severity and variability across intersections. Relative speed (in $\mu$) and relative distance (in $\sigma$) were also modeled as random effects (Fig. 5.1.1), although their 95% credible intervals overlapped zero. They were nevertheless retained on theoretical grounds, as their inclusion improved overall model fit. This specification reflects behavioral intuition: heterogeneity in speed effects is more pronounced in severity (location), whereas heterogeneity in clearance effects is more evident in variability (scale). All other covariates were treated as fixed effects, reflecting their comparatively stable behavioral and geometric influences. Fit statistics (Table 6) show modest but consistent improvements in DIC, WAIC, and LOOIC relative to the HBSFP specification, suggesting that the grouped random parameter structure adds explanatory nuance but produces smaller performance gains than in the V–V case.
@@
In the location parameter, several interpretable effects emerge. Relative distance is strongly negative and significant, confirming that reduced clearance to the roadway edge sharply lowers 2D-TTC and elevates severity. Relative deceleration also enters with a significant negative sign, contrasting with its positive effect in V–V near misses. This suggests that braking near infrastructure often reflects proactive yielding or evasive maneuvers, which mitigate severity rather than intensify it. Heading and steering differences are positive and significant, indicating that lateral deviations (e.g., turning toward or away from curbs) increase near-miss severity. Left-turn movement, by contrast, shows a significant negative association, suggesting that left-turn trajectories provide greater clearance to boundaries and reduce severity compared with right turns. This pattern may reflect geometric characteristics such as wider turning radii or reduced overlap with curbs and crosswalks during left-turn phasing. Finally, relative speed—although modeled as a random effect—does not reach significance in the pooled analysis, reinforcing that V–I severity is less sensitive to speed differentials than V–V encounters.
@@
The scale parameter results provide additional insight into variability. The negative random intercept indicates baseline differences across intersections, with some sites producing more predictable outcomes than others. Relative speed shows a positive and significant association with $\sigma$, implying that faster approaches to infrastructure increase variability in severity—in contrast to the V–V case, where speed compressed dispersion. This divergence reflects the diverse ways drivers respond to boundary proximity: some brake aggressively, while others maintain speed or adjust trajectory. Relative acceleration is also positive and significant, reinforcing that heterogeneous acceleration behavior near boundaries broadens outcome variability. By contrast, relative deceleration has a significant negative effect, indicating that decisive braking compresses the distribution into consistently critical outcomes. Jerk is similarly positive, confirming that abrupt throttle–brake oscillations amplify variability. Lateral dynamics play the opposite role: heading and steering differences both reduce variability, suggesting that once a lateral maneuver is committed, outcomes become more predictable within intersection geometry. Finally, left-turn movement increases variability despite lowering mean severity, consistent with heterogeneous layouts and driver strategies in left-turn contexts.
@@
The shape parameter remains negative and significant across intersections, supporting the Weibull domain for V–I near misses and confirming that these events have a bounded upper tail. This indicates a physical and behavioral ceiling on severity, constrained by roadway geometry and driver response capacity. Spatial variation in $\xi$ further underscores the influence of intersection design and infrastructure placement on the extremal properties of near-miss distributions. Taken together, the intersection-level results suggest that V–I near misses are shaped more by clearance, braking, and lateral maneuvers than by speed differentials. In contrast, V–V near misses are dominated by high speeds, which produce deterministic severity, whereas V–I outcomes exhibit greater variability driven by the interplay of infrastructure geometry, driver avoidance strategies, and intersection control.
5.1.2 Segment-level